OpenStudy (anonymous):

Solve M = 2x + 3y for y

OpenStudy (anonymous):

M=2x+3y Your goal should be to isolate y on one side

OpenStudy (anonymous):

y=13

OpenStudy (anonymous):

so it would be M - 2x = 2x + 3y -2x Subtract 2x from both sides of the equation M - 2x = 3y (M-2x)/3=3y/3 Divide both sides of the equation by 3 (M-2x)/3=y y=M/3-2x/3 or y=(M-2x)/3

OpenStudy (anonymous):

y=13?

OpenStudy (poofypenguin):

@Trexy answer seems very reasonable. I really don't think you can get a numerical value from this equations so y=13 seems out of question.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

lol Im sorry haha wrong problem disregard "y=13'

OpenStudy (anonymous):

so Im wrong sorry.. well okay.

OpenStudy (poofypenguin):

Wait! Is it actually? At a quick glance it looked fine to me! Let me have a closer look.

OpenStudy (poofypenguin):

@Trexy you're correct with the y=(M-2x)/3 solution.

OpenStudy (anonymous):

Ok.. Sorry..